OpenStudy (anonymous):

Identify the Exponent Law shown in the example below. (a2)3 = a2·3 = a6 Power of a Product Power of a Power Product of Powers None of the Above

OpenStudy (anonymous):

D

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

that's a power of a power.

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

\(\large (a^2)^3=a^2\cdot a^2\cdot a^2=a^{2+2+2}=a^6 \) so 2+2+2 = 3*2 =6